# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ B.2 Results using default hyperparameter configurations
Dataset FT-Transformer ResNet SAINT TabNet adult 0.893029 0.905838 0.870099 0.912781 analcatdata_authorship 0.999392 1.000000 0.999983 0.993186 analcatdata_dmft 0.553755 0.553675 0.526597 0.534271 balance-scale 0.988863 0.992229 0.991970 0.972816 bank-marketing 0.907667 0.926617 0.892316 0.927765 banknote-authentication 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.804580 0.843462 - 0.812061 blood-transfusion-service-center 0.713181 0.742088 0.723673 0.728919 breast-w 0.988615 0.991140 0.992220 0.984383 car 0.999758 0.998600 0.999828 0.931659 churn 0.915966 0.914732 0.910996 0.905642 climate-model-simulation-crashes 0.840724 0.904025 0.937306 0.825571 cmc 0.686016 0.687757 0.642394 0.689043 cnae-9 0.994801 0.996595 - 0.912423 connect-4 0.922969 0.926041 0.756318 0.856762 credit-approval 0.915482 0.916769 0.908623 0.875614 credit-g 0.731714 0.735071 0.744000 0.632571 cylinder-bands 0.908565 0.891759 0.909314 0.710240 diabetes 0.755846 0.789923 0.737127 0.785077 dna 0.988362 0.992218 0.520670 0.962713 dresses-sales 0.571921 0.536617 0.568144 0.560591 electricity 0.963347 0.930924 0.960991 0.911419 eucalyptus 0.917340 0.897582 0.904708 0.877684 first-order-theorem-proving 0.796282 0.793079 0.772449 0.743350 GesturePhaseSegmentationProcessed 0.827939 0.853272 0.893255 0.781506 har 0.999876 0.999859 - 0.999147 ilpd 0.724591 0.758030 0.713191 0.715948 Internet-Advertisements 0.973465 0.967077 - 0.892480 isolet 0.999463 0.999307 - 0.997706 jm1 0.723314 0.734238 0.652524 0.722615 jungle_chess_2pcs_raw_endgame_complete 0.998738 0.977410 0.999876 0.974173 kc1 0.804719 0.795200 0.742990 0.792858 kc2 0.805644 0.771497 0.742400 0.806986 kr-vs-kp 0.999792 0.999476 0.723052 0.987183 letter 0.999825 0.999864 0.999784 0.997271 madelon 0.770769 0.600713 - 0.559015 mfeat-factors 0.998765 0.998892 0.499849 0.993717 mfeat-fourier 0.977475 0.980419 0.971772 0.961111 mfeat-karhunen 0.997503 0.998097 0.998387 0.982592 mfeat-morphological 0.967733 0.969308 0.967478 0.963611 mfeat-pixel 0.997658 0.998676 0.553414 0.992500 mfeat-zernike 0.978039 0.980858 0.969257 0.966992 MiceProtein 1.000000 0.999963 1.000000 0.987043 nomao 0.992049 0.992530 0.499521 0.991441 numerai28.6 0.507813 0.517071 0.507780 0.522797 optdigits 0.999631 0.999837 0.999057 0.998476 ozone-level-8hr 0.893747 0.826296 0.881560 0.869228 pc1 0.852119 0.820008 0.866325 0.863233 pc3 0.810311 0.771759 0.804479 0.809443 pc4 0.944764 0.936765 0.931286 0.900752 pendigits 0.999740 0.999691 0.999785 0.999088 PhishingWebsites 0.996882 0.997134 0.996805 0.993856 phoneme 0.956543 0.938565 0.956949 0.933545 qsar-biodeg 0.916158 0.916804 0.918103 0.893489 satimage 0.992141 0.990613 0.985874 0.986280 segment 0.994709 0.993821 0.993989 0.992101 semeion 0.995507 0.996745 0.576269 0.957550 sick 0.997877 0.969015 0.991121 0.929353 spambase 0.983325 0.985056 0.981111 0.978240 splice 0.989898 0.990917 0.991932 0.972882 steel-plates-fault 0.959626 0.959356 0.948021 0.916561 texture 0.999976 0.999999 0.996944 0.999441 tic-tac-toe 0.998605 0.999375 0.996921 0.899715 vehicle 0.956404 0.963268 0.944376 0.923325 vowel 0.999618 0.999966 0.999888 0.986644 wall-robot-navigation 0.999757 0.998972 0.999104 0.997972 wdbc 0.994847 0.997080 0.997234 0.985323 wilt 0.994235 0.994057 0.988766 0.991840
Table 20 shows the raw results for MLP and RealMLP using the default hyperparameter configurations.
Table 20: Average test ROC-AUC per dataset for MLP and RealMLP using the default hyperparamater configurations across CV folds.
Dataset MLP RealMLP adult 0.897504 0.908495 analcatdata_authorship 0.999934 0.999952 analcatdata_dmft 0.554240 0.575077 balance-scale 0.995111 0.980107 bank-marketing 0.904699 0.816579 banknote-authentication 1.000000 1.000000 Bioresponse 0.560952 0.824996 blood-transfusion-service-center 0.762080 0.746119 breast-w 0.994222 0.993411 car 0.999678 1.000000 churn 0.903166 0.917012 climate-model-simulation-crashes 0.935571 0.947857 cmc 0.710134 0.700557 cnae-9 0.500463 0.992911 connect-4 0.915051 0.909829 credit-approval 0.931215 0.914531 credit-g 0.726875 0.758000 cylinder-bands 0.874205 0.906434 diabetes 0.829853 0.822211 dna 0.990128 0.988320 dresses-sales 0.536782 0.526601 electricity 0.950665 0.950555 eucalyptus 0.922173 0.903885 first-order-theorem-proving 0.782461 0.781809 GesturePhaseSegmentationProcessed 0.819054 0.890444 har 0.999848 0.999630 ilpd 0.748217 0.727899 Internet-Advertisements 0.982883 0.961953 isolet 0.847095 0.999135 jm1 0.726646 0.721977 jungle_chess_2pcs_raw_endgame_complete 0.998486 0.996257 kc1 0.801565 0.806604 kc2 0.840419 0.829826 kr-vs-kp 0.999765 0.998737 letter 0.999640 0.999820 madelon 0.500000 0.915592 mfeat-factors 0.998668 0.999075 mfeat-fourier 0.978653 0.974028 mfeat-karhunen 0.998582 0.999439 mfeat-morphological 0.965494 0.968706 mfeat-pixel 0.946632 0.999500 mfeat-zernike 0.980681 0.965872 MiceProtein 0.999963 1.000000 nomao 0.991436 0.983015 numerai28.6 0.513601 0.522412 optdigits 0.999454 0.999927 ozone-level-8hr 0.906572 0.822254 pc1 0.853077 0.828996 pc3 0.784672 0.768438 pc4 0.940799 0.906347 pendigits 0.999687 0.999850 PhishingWebsites 0.996479 0.994417 phoneme 0.948168 0.952913 qsar-biodeg 0.924529 0.911174 satimage 0.990975 0.986944 segment 0.993795 0.994189 semeion 0.968306 0.998289 sick 0.989590 0.976784 spambase 0.983168 0.978382 splice 0.990919 0.991318 steel-plates-fault 0.963250 0.955133 texture 0.999956 0.999992 tic-tac-toe 0.999145 0.997548 vehicle 0.944588 0.961117 vowel 0.997520 0.999641 wall-robot-navigation 0.999245 0.998582 wdbc 0.994219 0.998021 wilt 0.994105 0.993080
@@
Table 21: Average test ROC-AUC per dataset for meta-learned neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by "-".
@@ Appendix C Datasets
Appendix D Hyperparameter analysis
In this section, we analyze the impact of individual hyperparameters on the performance metric. The x-axis represents the hyperparameters, while the y-axis denotes the ROC-AUC performance. These plots provide an overview of the performance landscape for each hyperparameter, illustrating their influence on model effectiveness.
D.1 CatBoost
D.2 ResNet
D.3 MLP-PLR
D.4 RealMLP
D.5 XGBoost
D.6 FT-Transformer
D.7 SAINT
D.8 TabNet
D.9 XTab
D.10 CARTE
D.11 TP-BERTa
Appendix E Analysis of Dataset Characteristics: Instances and Features
@@
E.1 Datasets with fewer than 1000 instances
@@
In this section, we focus on datasets with fewer than 100 features and fewer than 1000 instances, resulting in a total of 18 datasets used in our study. Consequently, most methods in Figure 27 are evaluated on 18 datasets. However, there are a few exceptions: TabPFN is incompatible with one dataset, "vowel," due to it containing more than 10 classes; XTab excludes 2 datasets that were part of its pretraining phase; and TP-Berta encounters memory limitations on 10 out of the 18 datasets, reducing its coverage.
@@
Figure 27: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
@@
Figure 27 reveals that AutoGluon achieves the strongest overall performance, closely followed by TabPFN. Among feedforward networks, MLP and RealMLP both rank well, though MLP shows a tighter (i.e. more robust) interquartile range. Among the other dataset-specific neural networks, FT-Transformer and SAINT perform comparably. Interestingly, MLP-like methods also show a lower median rank than the classical CatBoost and XGBoost, although CatBoost occasionally achieves ranks as low as 2.5. By contrast, TabNet and the fine-tuning–based models generally exhibit the weakest performance.
@@
Figure 28: Left: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range. Right: Comparative analysis of all the methods.
@@
Similarly, Figure 28 shows a boxplot on the left—evaluated on the same datasets but excluding TP-BERTa—and a critical difference diagram on the right. A clear pattern emerges: in the small-data domain, dataset-specific neural architectures (e.g., MLP with PLR embeddings, RealMLP and FT-Transformer) display highly competitive performance, surpassing even CatBoost and XGBoost.
E.2 Datasets with 1,000 to 5,000 instances
@@
Following the previous analysis, we now focus on datasets with 1,000 to 5,000 instances and fewer than 100 features. Figure 29 illustrates the results for this subset of datasets. Similar to the small data domain, AutoGluon remains the top-performing method. However, an interesting shift occurs, CatBoost shows a significant improvement in performance, achieving the second-best overall rank, while XGBoost maintains a performance level similar to the smaller datasets. Additionally, dataset-specific neural networks continue to outperform meta-learned neural networks, with the MLP with PLR embeddings standing out due to its strong performance. It exhibits a better median rank and a narrower interquartile range compared to XGBoost.
@@
Figure 29: Distribution of ranks for all the methods in the datasets withh 1000 to 5000 instances, and less than 100 features. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
@@
In Figure 30, we exclude TP-BERTa again to ensure a reasonable number of common datasets, resulting in a total of 19 datasets. The left plot tells a similar story, with XGBoost now achieving the same median rank as the MLP with PLR embeddings. The right plot presents a critical difference diagram, showing AutoGluon, CatBoost, MLP, XGBoost and RealMLP as the top-performing methods. Among them, AutoGluon is statistically significantly better than all other methods, except for the aforementioned top performers.
@@
Table 26: Classifier Performance for Instance Range: 1000-5000 and Feature Range: >1000
@@
E.3 Datasets with 5,000 to 10,000 instances
Table 27: Classifier Performance for Instance Range: 5000-10000 and Feature Range: <=100
Dataset GPhaseSeg first-ord-TP optdigits phoneme satimage texture wall-rob-nav
AutoGluon 0.936667 0.835425 0.999925 0.973342 0.993557 0.999998 0.999993
CARTE 0.798024 0.764092 0.999112 0.948702 0.988038 0.999541 0.999505
CatBoost 0.916674 0.831775 0.999844 0.968024 0.991978 0.999948 0.999990
MLP 0.911434 0.798811 0.999794 0.967617 0.992308 0.999990 0.999689
RealMLP 0.901441 0.795637 0.999968 0.966456 0.993034 0.999999 0.998720
FT-Transformer 0.895166 0.796707 0.999616 0.965071 0.993516 0.999983 0.999900
XGBoost 0.916761 0.834883 0.999855 0.967421 0.992114 0.999940 0.999981

### v3

@@ D.2 Results using default hyperparameter configurations
Dataset FT-Transformer MLP RealMLP ResNet SAINT TabM TabNet adult 0.893029 0.897504 0.909085 0.905838 0.870099 0.908670 0.912781 analcatdata_authorship 0.999392 0.999934 0.999952 1.000000 0.999983 1.000000 0.993186 analcatdata_dmft 0.553755 0.554240 0.575007 0.553675 0.526597 0.539602 0.534271 balance-scale 0.988863 0.995111 0.980107 0.992229 0.991970 0.998859 0.972816 bank-marketing 0.907667 0.904699 0.814657 0.926617 0.892316 0.931979 0.927765 banknote-authentication 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.804580 0.560952 0.824996 0.843462 - 0.872512 0.812061 blood-transfusion-service-center 0.713181 0.762080 0.746119 0.742088 0.723673 0.727600 0.728919 breast-w 0.988615 0.994222 0.993411 0.991140 0.992220 0.995038 0.984383 car 0.999758 0.999678 1.000000 0.998600 0.999828 1.000000 0.931659 churn 0.915966 0.903166 0.917916 0.914732 0.910996 0.927037 0.905642 climate-model-simulation-crashes 0.840724 0.935571 0.947857 0.904025 0.937306 0.946051 0.825571 cmc 0.686016 0.710134 0.700557 0.687757 0.642394 0.692772 0.689043 cnae-9 0.994801 0.500463 0.992911 0.996595 - 0.997415 0.912423 connect-4 0.922969 0.915051 0.909829 0.926041 0.756318 0.938629 0.856762 credit-approval 0.915482 0.931215 0.914193 0.916769 0.908623 0.920201 0.875614 credit-g 0.731714 0.726875 0.758571 0.735071 0.744000 0.782714 0.632571 cylinder-bands 0.908565 0.874205 0.904906 0.891759 0.909314 0.924863 0.710240 diabetes 0.755846 0.829853 0.822211 0.789923 0.737127 0.789142 0.785077 dna 0.988362 0.990128 0.988320 0.992218 0.520670 0.993741 0.962713 dresses-sales 0.571921 0.536782 0.525944 0.536617 0.568144 0.542365 0.560591 electricity 0.963347 0.950665 0.950555 0.930924 0.960991 0.959880 0.911419 eucalyptus 0.917340 0.922173 0.903412 0.897582 0.904708 0.925170 0.877684 first-order-theorem-proving 0.796282 0.782461 0.781809 0.793079 0.772449 0.814144 0.743350 GesturePhaseSegmentationProcessed 0.827939 0.819054 0.890444 0.853272 0.893255 0.874995 0.781506 har 0.999876 0.999848 0.999630 0.999859 - 0.999937 0.999147 ilpd 0.724591 0.748217 0.727899 0.758030 0.713191 0.759658 0.715948 Internet-Advertisements 0.973465 0.982883 0.961953 0.967077 - 0.981634 0.892480 isolet 0.999463 0.847095 0.999135 0.999307 - 0.999671 0.997706 jm1 0.723314 0.726646 0.721977 0.734238 0.652524 0.738839 0.722615 jungle_chess_2pcs_raw_endgame_complete 0.998738 0.998486 0.996257 0.977410 0.999876 0.998544 0.974173 kc1 0.804719 0.801565 0.806604 0.795200 0.742990 0.820436 0.792858 kc2 0.805644 0.840419 0.829826 0.771497 0.742400 0.818842 0.806986 kr-vs-kp 0.999792 0.999765 0.998737 0.999476 0.723052 0.999796 0.987183 letter 0.999825 0.999640 0.999820 0.999864 0.999784 0.999918 0.997271 madelon 0.770769 0.500000 0.915592 0.600713 - 0.758018 0.559015 mfeat-factors 0.998765 0.998668 0.999075 0.998892 0.499849 0.999589 0.993717 mfeat-fourier 0.977475 0.978653 0.974028 0.980419 0.971772 0.986450 0.961111 mfeat-karhunen 0.997503 0.998582 0.999439 0.998097 0.998387 0.998958 0.982592 mfeat-morphological 0.967733 0.965494 0.968706 0.969308 0.967478 0.967408 0.963611 mfeat-pixel 0.997658 0.946632 0.999500 0.998676 0.553414 0.999317 0.992500 mfeat-zernike 0.978039 0.980681 0.965872 0.980858 0.969257 0.978049 0.966992 MiceProtein 1.000000 0.999963 1.000000 0.999963 1.000000 0.999991 0.987043 nomao 0.992049 0.991436 0.983015 0.992530 0.499521 0.994784 0.991441 numerai28.6 0.507813 0.513601 0.522412 0.517071 0.507780 0.523854 0.522797 optdigits 0.999631 0.999454 0.999927 0.999837 0.999057 0.999927 0.998476 ozone-level-8hr 0.893747 0.906572 0.822254 0.826296 0.881560 0.921495 0.869228 pc1 0.852119 0.853077 0.828996 0.820008 0.866325 0.891588 0.863233 pc3 0.810311 0.784672 0.768438 0.771759 0.804479 0.832408 0.809443 pc4 0.944764 0.940799 0.906347 0.936765 0.931286 0.949946 0.900752 pendigits 0.999740 0.999687 0.999850 0.999691 0.999785 0.999748 0.999088 PhishingWebsites 0.996882 0.996479 0.994417 0.997134 0.996805 0.997615 0.993856 phoneme 0.956543 0.948168 0.952913 0.938565 0.956949 0.963998 0.933545 qsar-biodeg 0.916158 0.924529 0.911174 0.916804 0.918103 0.934609 0.893489 satimage 0.992141 0.990975 0.986944 0.990613 0.985874 0.993552 0.986280 segment 0.994709 0.993795 0.994189 0.993821 0.993989 0.994829 0.992101 semeion 0.995507 0.968306 0.998289 0.996745 0.576269 0.998174 0.957550 sick 0.997877 0.989590 0.976784 0.969015 0.991121 0.995715 0.929353 spambase 0.983325 0.983168 0.978382 0.985056 0.981111 0.988298 0.978240 splice 0.989898 0.990919 0.991318 0.990917 0.991932 0.995071 0.972882 steel-plates-fault 0.959626 0.963250 0.955133 0.959356 0.948021 0.966012 0.916561 texture 0.999976 0.999956 0.999992 0.999999 0.996944 0.999997 0.999441 tic-tac-toe 0.998605 0.999145 0.997548 0.999375 0.996921 0.999904 0.899715 vehicle 0.956404 0.944588 0.961117 0.963268 0.944376 0.962549 0.923325 vowel 0.999618 0.997520 0.999641 0.999966 0.999888 0.999854 0.986644 wall-robot-navigation 0.999757 0.999245 0.998582 0.998972 0.999104 0.999520 0.997972 wdbc 0.994847 0.994219 0.998021 0.997080 0.997234 0.997765 0.985323 wilt 0.994235 0.994105 0.993080 0.994057 0.988766 0.994455 0.991840
@@
Table 23: Average test ROC-AUC per dataset for meta-learned neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by ”-”.
@@ Appendix E Datasets
We evaluate all methods on the OpenML CC-18 benchmark. Four very large image datasets are excluded a priori due to memory issues affecting most baselines: Devnagari-Script (OpenML ID 40923), MNIST (554), Fashion-MNIST (40996), and CIFAR-10 (40927). This leaves 68 datasets as the default pool. Some methods are run on fewer datasets due to memory limits, method-specific constraints, or pretraining overlap. Table 26 reports the coverage per method and the reason for any missing datasets.
Table 26: Dataset coverage on OpenML CC-18 after excluding four image datasets (IDs: 40923, 554, 40996, 40927). Default pool = 68 datasets.
Method Used Missing Primary reason for missing datasets AutoGluon 68 0 CatBoost 68 0 LightGBM 68 0 XGBoost 68 0 RealMLP 68 0 TabM 68 0 ResNet 68 0 FT-Transformer 68 0 MLP 67 1 Memory constraint on one dataset SAINT 61 7 Memory constraints TabNet 62 6 Memory constraints CARTE 62 6 Memory constraints TabICL 68 0 TP-BERTa 15 53 Memory constraints TabPFN 17 51 Method limits: $\leq$1000 samples, $\leq$100 features, $\leq$10 classes TabPFNv2 49 19 Method limits: $\leq$10,000 samples, $\leq$500 features, $\leq$10 classes XTab 55 13 Excluded due to pretraining overlap
Appendix F Comparison with AutoML methods
Figure 27: Distribution of ranks for the Deep Learning Models (13 methods) and AutoML (1 method) classifier families. Left: AutoGluon with hyperparameter optimization (HPO). Right: AutoGluon in its recommended configuration. The boxplot illustrates the rank spread, with medians represented by red lines and whiskers showing the range.
Figure 27 presents boxplots of the rank distributions for all Deep Learning methods compared to AutoGluon. The left-hand side shows results against the HPO-tuned version of AutoGluon. TabPFNv2 and TabICL achieve the best overall performance, sharing the same mean rank, with TabICL having a slightly better median rank. Both outperform all other methods, including AutoGluon (HPO). TabM also performs competitively with a median rank around 3, while most other Deep Learning methods rank lower.
The right-hand side compares against the recommended version of AutoGluon and reveals a similar picture. TabPFNv2 matches AutoGluon in both mean and median rank, though AutoGluon’s interquartile range extends slightly lower, indicating marginally better performance. TabICL attains the same mean rank as TabPFNv2 and AutoGluon, but with a better median rank.
Consistent with our previous analyses, we also present critical difference (CD) diagrams to summarize the ranking comparisons. Figure 28 shows, on the left, the CD diagram comparing AutoGluon with HPO against the Deep Learning methods across all datasets, and on the right, the diagram for AutoGluon with its recommended configuration.
On the left, TabICL and TabPFNv2 achieve the best average rank of 3, followed by TabM and then AutoGluon (HPO). The diagram indicates that differences among TabICL, TabPFNv2, and TabM are not statistically significant; however, TabICL and TabPFNv2 significantly outperform all remaining methods.
On the right, TabICL is the top-ranked method, followed closely by TabPFNv2 and AutoGluon. The absence of connecting bars for these three methods indicates that their performance is statistically significantly better than that of all other Deep Learning baselines.
We further perform the same analysis in the small-data setting, defined as datasets with $\leq 1000$ instances, which allows us to include TabPFN in the comparison. The results are shown in Figure 29.
In the left diagram, TabPFNv2 and TabICL lead the rankings, closely followed by TabM, all outperforming the HPO-tuned version of AutoGluon. In contrast, the right diagram shows that AutoGluon with its recommended configuration ranks highest, surpassing even TabICL and TabPFNv2, although the differences are not statistically significant.
Appendix G Analysis of Dataset Characteristics: Instances and Features
@@
G.1 Datasets with fewer than 1000 instances
@@
In this section, we focus on datasets with fewer than 100 features and fewer than 1000 instances, resulting in a total of 18 datasets used in our study. Consequently, most methods in Figure 30 are evaluated on 18 datasets. However, there are a few exceptions: TabPFN and TabPFNv2 are incompatible with one dataset, ”vowel,” due to it containing more than 10 classes; XTab excludes 2 datasets that were part of its pretraining phase; and TP-Berta encounters memory limitations on 10 out of the 18 datasets, reducing its coverage.
@@
Figure 30: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range.
@@
Figure 30 reveals that AutoGluon achieves the strongest overall performance, closely followed by TabPFNv2 and TabICL. Among feedforward networks, TabM, MLP and RealMLP rank well, though TabM reaches a lower rank, while RealMLP has a better mean and median rank compared to MLP. Among the other dataset-specific neural networks, FT-Transformer and SAINT perform comparably. Interestingly, MLP-like methods also show a lower median rank than the classical CatBoost, LightGBM and XGBoost, although CatBoost occasionally achieves ranks as low as 2.5. By contrast, TabNet and the fine-tuning–based models generally exhibit the weakest performance.
@@
Figure 31: Left: Distribution of ranks for all the methods in the small data domain. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range. Right: Comparative analysis of all the methods.
@@
Similarly, Figure 31 shows a boxplot on the left, evaluated on the same datasets but excluding TP-BERTa—and a critical difference diagram on the right. A clear pattern emerges: in the small-data domain, in-context learning methods, are highly competitive, followed by dataset-specific neural architectures (e.g., TabM, MLP with PLR embeddings and RealMLP), surpassing even CatBoost, LightGBM and XGBoost.
G.2 Datasets with 1,000 to 5,000 instances
@@
Following the previous analysis, we now examine datasets with 1,000–5,000 instances and fewer than 100 features. The results are shown in Figure 32. Similar to the small-data setting, TabICL and TabPFNv2 dominate, outperforming even AutoGluon. A notable shift, however, is the strong performance of CatBoost, which rises to fourth place overall, just behind TabM in third, while XGBoost maintains performance levels comparable to those in the smaller datasets. Furthermore, dataset-specific neural networks continue to outperform fine-tuned networks, with TabM and MLP with PLR embeddings standing out for their strong performance. Both achieve better median ranks and narrower interquartile ranges than XGBoost and LightGBM.
@@
Figure 32: Distribution of ranks for all the methods in the datasets with 1000 to 5000 instances, and less than 100 features. The boxplot illustrates the rank spread, with medians represented by black lines, means represented by diamonds and whiskers showing the range.
@@
In Figure 33, we exclude TP-BERTa again to ensure a reasonable number of common datasets, resulting in a total of 19 datasets. The left plot tells a similar story, with XGBoost now achieving slightly better median rank as the MLP with PLR embeddings. The right plot presents a critical difference diagram, showing TabICL, TabPFNv2, and AutoGluon as the top-performing methods.
@@
Table 29: Classifier Performance for Instance Range: 1000-5000 and Feature Range: >> 1000
@@
G.3 Datasets with 5,000 to 10,000 instances
Table 30: Classifier Performance for Instance Range: 5000-10000 and Feature Range: $\leq$ 100
Dataset GPhaseSeg first-ord-TP optdigits phoneme satimage texture wall-rob-nav AutoGluon 0.936667 0.835425 0.999925 0.973342 0.993557 0.999998 0.999993 CARTE 0.798024 0.764092 0.999112 0.948702 0.988038 0.999541 0.999505 CatBoost 0.916674 0.831775 0.999844 0.968024 0.991978 0.999948 0.999990 FT-Transformer 0.895166 0.796707 0.999616 0.965071 0.993516 0.999983 0.999900 LightGBM 0.920034 0.833949 0.999818 0.965147 0.991309 0.999890 0.999968 MLP 0.911434 0.798812 0.999794 0.967617 0.992308 0.999991 0.999689 RealMLP 0.901441 0.795637 0.999968 0.966456 0.993034 0.999999 0.998720 ResNet 0.914196 0.784636 0.999927 0.963591 0.991995 0.999999 0.999042 SAINT 0.919006 0.802392 0.999841 0.960382 0.992630 0.999976 0.999844 TabPFNv2 0.936548 0.825502 0.999897 0.973546 0.995122 0.999963 0.999936 TabM 0.933828 0.818255 0.999939 0.971200 0.994291 0.999997 0.999912 TabNet 0.850596 0.774094 0.998871 0.956279 0.987482 0.999763 0.997585 TabICL 0.951408 0.834629 0.999989 0.977493 0.993373 1.000000 0.999610 XGBoost 0.916761 0.834883 0.999855 0.967421 0.992114 0.999940 0.999981 XTab 0.886960 0.798803 0.999712 0.961749 0.992918 0.999962 0.999846